OpenStudy (anonymous):

Please Help Exponential Decay Word problem suppose a person takes a fever reducer medicine that contains ibuprofen. After two hours, the amount of ibuprofen remained in the body is 120 mg and after 6 hours the amount reduced to 30 mg. A. find an exponential decay function that models the amount of Ibuprofen in the body after t hours.

OpenStudy (anonymous):

120mg 30mg 2h 6h decreased by 1/4 in 4h so \[\frac{1}{4} = 120e^{-k(4)} = [A]_{t} = [A]_{0}e^{-kt}\] solve for -k

OpenStudy (anonymous):

Ignore that 120 in the first part you know that it took 4 hours to get to 30mg so solve for k by \[\frac{1}{4} = e^{-k(4)}\] e is the natural log